# Transition Metal Dichalcogenides Multijunction Solar Cells Toward the Multicolor Limit

### v1

@@ 4 Results
Luminescent coupling and a power-budget view of entropy vs thermalization. Figure 9 quantifies when luminescent coupling becomes appreciable in reciprocal stacks and provides a power-budget decomposition for the representative $N=5$ ladder. Panel (a) shows that the downward coupling current $J_{\mathrm{LC}}$ at the system MPP increases strongly with concentration and is amplified by high ERE via higher operating voltage. Panel (b) shows that the coupling ratio grows rapidly when adjacent bandgaps become closely spaced (small $\Delta E_{g}$), a condition naturally approached in dense high-$N$ ladders. Panel (c) provides a compact power budget for the representative $N=5$ ladder: the upward luminescence power is plotted as an entropy-loss proxy, while coupling-induced thermalization is plotted separately. This decomposition clarifies the key design tension: suppressing upward emission can reduce entropy loss (and increase voltage), but downward emission into lower junctions can still create thermalization loss if energy is not efficiently converted.
@@
Importantly, coupling is not universally detrimental. When the lower cell converts the coupled photons efficiently and the bandgap spacing is sufficiently large, downward luminescence can partially compensate current deficits and relax the need for perfect spectral splitting. However, in dense ladders with small bandgap spacing, coupled photons arrive with excess energy relative to the lower bandgap and are more likely to be dissipated as heat, effectively reintroducing a thermalization channel that grows with concentration and ERE. For the representative $N=5$ ladder at full concentration, the reciprocal ideal-optics stack loses $\sim 1.24\%$ of the incident solar power as upward-emitted luminescence, whereas in the idealized nonreciprocal model this upward channel is eliminated and the electrical output increases by $\sim 1.5\%$ absolute, with part of the redirected luminescence appearing as coupling-induced thermalization (Figure 9c).
@@
Reciprocal vs nonreciprocal headroom (conservative “plug-in” comparison). Nonreciprocal multijunction photovoltaics has been proposed as a route to approach the Landsberg limit by suppressing entropy losses associated with reciprocal emission pathways.13 However, for single-junction solar cells, nonreciprocity does not increase the SQ efficiency limit under detailed balance.12 To place these insights in the present context, we benchmark the potential headroom by applying the idealized nonreciprocal multijunction model to the same bandgap ladders optimized under reciprocity (Option A, conservative comparison).
@@
Figure 10 compares full-concentration efficiencies versus $N$ for reciprocal and nonreciprocal cases, for both unconstrained bandgaps and the conservative TMD window. In the TMD window, the reciprocal efficiency plateaus near $\sim 63.4\%$ at large $N$, whereas the nonreciprocal model rises to $\sim 67.7\%$ at $N=50$ (Figure 10a). Figure 10b shows the central physical signature: the nonreciprocal model suppresses upward-emitted luminescence power (entropy-loss proxy), whereas the reciprocal case must emit upward to maintain detailed balance. The filled markers in Figure 10a further show that reciprocal stacks without intermediate mirrors incur additional penalties from two-sided emission and coupling, highlighting the practical importance of intermediate mirrors or angular filters in realistic reciprocal designs.
In the conservative Option A comparison, the nonreciprocal model is applied to the reciprocity-optimized ladder without re-optimizing bandgaps, so the improvement should be interpreted as a lower bound on the possible headroom. In practice, achieving the ideal nonreciprocal behavior is nontrivial, but the comparison clarifies what must be suppressed in reciprocal stacks: upward emission that carries entropy, and downward emission that drives coupling thermalization. Intermediate mirrors, angular filters, and spectral-selective contacts can address part of this gap without requiring fully nonreciprocal photonics.
@@
Figure 10: Reciprocal vs nonreciprocal headroom under full concentration. (a) Efficiency versus junction number for reciprocal split-spectrum multijunctions (one-sided emission, no coupling), and for the idealized nonreciprocal multijunction model of Fan et al.13 applied to the same reciprocity-optimized bandgap ladders (Option A). Filled markers show a reciprocal-stack estimate without intermediate mirrors (two-sided emission + downward luminescent coupling) for $N\leq 10$. (b) Upward-emitted luminescence power fraction $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy. In the idealized nonreciprocal model used here, the upward-emitted luminescence is identically zero, so the nonreciprocal series in (b) are shown with a slight horizontal offset for visibility. Horizontal lines indicate literature multicolor and Landsberg limits for comparison (values depend weakly on $T_{s}$).13
5 Discussion: design rules for transfer-printed vdW/TMD multijunction photovoltaics
We summarize design rules suggested by the above results.
@@ 5 Discussion: design rules for transfer-printed vdW/TMD multijunction photovoltaics
The DP optimizer provides optimal ladders and reveals the large-$N$ plateau imposed by a constrained bandgap window. Within the conservative 1.0–2.1 eV TMD window, $N\approx 5$ already approaches the plateau, suggesting that the most immediate experimental objective is to realize high-quality $N\leq 5$ stacks with excellent optics rather than extremely large $N$.
@@
Voltage penalties from finite ERE accumulate across subcells (Figure 6). High-NN stacks therefore demand high radiative quality in every junction, reinforcing the importance of passivation, low-loss contacts, and photon extraction.11, 7
(3) Control of emission channels and luminescent coupling is essential.
Without intermediate mirrors, two-sided emission introduces an entropy penalty and enables downward luminescent coupling (Figure 9a–b). Coupling can be beneficial for current in lower cells but also introduces thermalization losses and can reduce the emitter voltage. Practical multijunction architectures will likely require selective optical elements (intermediate mirrors, angular filters, or photonic cavities) designed to maximize useful coupling while minimizing entropy penalties.15, 19, 5
@@
Ultrathin absorbers cannot be treated as step absorbers; finite absorptance reduces both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$ through reciprocity (Figure 7). Nanophotonic bounds and thickness constraints limit the achievable broadband path-length enhancement (Figure 8). These effects motivate moderate absorber thicknesses (tens to hundreds of nm) and co-designed photonic structures rather than extreme monolayer limits for high-efficiency multijunction conversion.
@@
Consistent with Fan et al., nonreciprocity does not improve the SQ limit for single-junction devices.12 In multijunction architectures, however, reshaping emission pathways can reduce entropy losses and provide headroom toward Landsberg-like limits.13 Our conservative plug-in comparison (Figure 10) quantifies this headroom for TMD-optimized ladders while preserving the paper’s main focus: reciprocal multijunction design and realistic constraints.
@@ 6 Conclusions
We developed an unlimited-junction detailed-balance framework tailored to transfer-printed vdW/TMD multijunction photovoltaics and used it to compute optimal bandgap ladders under unconstrained and conservative TMD bandgap windows. A realistic bandgap window produces a large-$N$ efficiency plateau, motivating $N\sim 5$ as a practical near-term target for TMD-based stacks. We incorporated luminescence thermodynamics and reciprocity to quantify ERE penalties, emission-channel entropy, and luminescent coupling, and we introduced a power-budget view in which upward-emitted luminescence serves as a directly computable entropy-loss proxy while coupling thermalization is tracked separately. Finally, we benchmarked the remaining thermodynamic headroom by applying an idealized nonreciprocal multijunction model to our reciprocally optimized ladders, showing negligible benefit for single junctions but substantial headroom for multijunction stacks. These results provide quantitative design rules for approaching multicolor efficiencies with transfer-printed vdW multijunction photovoltaics.
@@
The numerical results and figures in this work are reproducible from the Python implementation of the detailed-balance, luminescent-coupling, and optimization models described in the main text and Supporting Information. To keep the manuscript self-contained for submission, the Supporting Information includes the full derivations, parameter tables, and additional figures. The Python source code and the raw numerical data required to regenerate every figure and table are available from the authors upon reasonable request and will be deposited in a public repository upon publication.
Supporting Information
The Supporting Information (SI) is provided as a separate PDF file. The SI provides detailed derivations, parameter tables, and additional figures that support the main text: (i) Supporting Information, Section 1 (Constants, geometry, and spectral model): solar/blackbody spectral definitions and constants; (ii) Supporting Information, Section 2 (Single-junction detailed balance): full detailed-balance derivations, including the Lambert-$W$ maximum-power condition; (iii) Supporting Information, Section 3 (Unlimited-junction split-spectrum multijunction model): the dynamic-programming (DP) optimizer used to compute optimal bandgap ladders for arbitrary $N$ and its discretization/convergence checks; (iv) Supporting Information, Section 4 (Reciprocity, emission channels, and the role of mirrors): reciprocity-based voltage-loss and entropy-loss decompositions, including the role of external radiative efficiency (ERE) and optical outcoupling; (v) Supporting Information, Section 5 (Luminescent coupling chain model): the luminescent-coupling chain model, together with the definition and computation of the upward emitted luminescence power and a coupling thermalization/heat proxy; (vi) Supporting Information, Section 6 (TMD optical response and excitonic absorptance model): a TMD material database (representative bandgaps, excitonic absorption spectra, and representative $n(\lambda)$$k(\lambda)$ dispersion conventions) and mapping of candidate stacks to the representative $N=5$ target ladder; (vii) Supporting Information, Section 7 (Nanophotonic bounds and Miller thickness constraints): nanophotonic bounds and thickness constraints relevant to ultrathin absorbers; (viii) Supporting Information, Section 8 (Nonreciprocal multijunction benchmark); (ix) Supporting Information, Section 9 (Extended bandgap-ladder tables for reproducibility); (x) Supporting Information, Section 10 (Additional figures for the representative $N=5$ stack; Figures S1 and S2); (xi) Supporting Information, Section 11 (Full ladder lists in text form).

### v3

@@ 4 Results
Luminescent coupling and a power-budget view of entropy vs thermalization. Figure 9 quantifies when luminescent coupling becomes appreciable in reciprocal stacks and provides a power-budget decomposition for the representative $N=5$ ladder. Figure 9(a) shows that the downward coupling current $J_{\mathrm{LC}}$ at the system MPP increases strongly with concentration and is amplified at high ERE due to higher operating voltage. Figure 9(b) shows that the coupling ratio grows rapidly, when adjacent bandgaps become closely spaced (small $\Delta E_{g}$) (i.e., a condition that is naturally approached in dense high-$N$ ladders). Figure 9(c) provides a compact power budget for the representative $N=5$ ladder: the upward luminescence power is plotted as an entropy-loss proxy, while coupling-induced thermalization is plotted separately. This decomposition clarifies the key design point: suppressing upward emission can reduce entropy loss (and increase voltage), but downward emission into lower junctions can still lead to thermalization loss, if energy is not efficiently converted.
@@
Importantly, coupling is not universally detrimental. When the lower cell converts the coupled photons efficiently and the bandgap spacing is sufficiently large, downward luminescence can partially compensate current deficits and relax the need for perfect spectral splitting. However, in dense ladders with small bandgap spacing, coupled photons arrive with excess energy relative to the lower bandgap. As a result, these coupled photons are more likely to be dissipated as heat, effectively reintroducing a thermalization channel that grows with concentration and ERE. For the representative $N=5$ ladder at full concentration, the reciprocal ideal-optics stack loses $\sim 1.24\%$ of the incident solar power as upward-emitted luminescence, whereas in the idealized nonreciprocal model, this upward channel is eliminated and the electrical output increases by $\sim 1.5\%$ absolute, with part of the redirected luminescence, which appear as coupling-induced thermalization (Figure 9c).
@@
Reciprocal vs nonreciprocal efficiency gain. Nonreciprocal multijunction photovoltaics has been proposed as a route to approach the Landsberg limit by suppressing entropy losses, which are associated with reciprocal emission pathways.[11] However, for single-junction solar cells, nonreciprocity does not increase the SQ efficiency limit under detailed balance.[10] To place these insights in the present context, we benchmark the potential efficiency gain by applying the idealized nonreciprocal multijunction model to the same bandgap ladders, that are optimized under reciprocity (Option A, conservative comparison).
@@
Figure 10 compares full-concentration efficiencies versus $N$ for reciprocal and nonreciprocal cases, for both unconstrained bandgaps and the conservative TMD window. In the TMD window, the reciprocal efficiency saturates near $\sim 63.4\%$ at large $N$, whereas the nonreciprocal model reaches to $\sim 67.7\%$ at $N=50$ (Figure 10a): extended bandgap ladder-tables are available in Tables S7–S10, Supplement Information. Figure 10b shows the critical physical signature: the nonreciprocal model suppresses upward-emitted luminescence power (entropy-loss proxy), whereas the reciprocal case must emit upward to maintain detailed balance. The filled markers in Figure 10a further show that reciprocal stacks without intermediate mirrors incur additional penalties from two-sided emission and coupling, highlighting the practical importance of intermediate mirrors or angular filters[4] in realistic reciprocal designs. In the conservative Option A comparison, the nonreciprocal model is implemented to the reciprocity-optimized ladder without re-optimizing bandgaps, so the improvement should be interpreted as a lower bound on the possible efficiency gain. In practice, achieving the ideal nonreciprocal behavior is nontrivial, but the comparison clarifies what must be suppressed in reciprocal stacks: upward emission that carries entropy and downward emission that drives coupling thermalization. Intermediate mirrors, angular filters, and spectral-selective contacts[4] can address part of this gap without requiring fully nonreciprocal photonics.
@@
Figure 10: Reciprocal vs nonreciprocal efficiency gain under full concentration. (a) Efficiency versus junction number for reciprocal split-spectrum multijunctions (one-sided emission, no coupling), and for the idealized nonreciprocal multijunction model of Fan et al.[11] applied to the same reciprocity-optimized bandgap ladders (Option A). Filled markers show a reciprocal-stack estimate without intermediate mirrors (two-sided emission + downward luminescent coupling) for $N\leq 10$. (b) Upward-emitted luminescence power fraction $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy. In the idealized nonreciprocal model used here, the upward-emitted luminescence is identically zero, so the nonreciprocal series in (b) are shown with a slight horizontal offset for visibility. Horizontal lines indicate literature multicolor and Landsberg limits for comparison (values depend weakly on $T_{s}$).[11]
5 Discussion: design rules for TMD multijunction cells
@@
The DP optimizer provides optimal ladders and reveals the large-$N$ plateau imposed by a constrained bandgap window. Within the conservative 1.0–2.1 eV TMD window, $N\approx 5$ already approaches the efficiency plateau, suggesting that the most immediate experimental objective is to realize high-quality $N\leq 5$ stacks with excellent optics rather than extremely large $N$.
@@
Voltage penalties from finite ERE accumulate across subcells (Figure 6). High-NN stacks, therefore, demand high radiative quality in every junction, highlighting the importance of passivation, low-loss contacts, and photon extraction.[9, 6]
(3) Control of emission channels and luminescent coupling are essential.
Without intermediate mirrors, two-sided emission introduces an entropy penalty and enables downward luminescent coupling (Figure 9a–b). Coupling can be beneficial for current in lower cells. However, it also introduces thermalization losses and reduces the voltage. Practical multijunction architectures will likely require selective optical elements (intermediate mirrors, angular filters, or photonic cavities) designed to maximize useful coupling while minimizing entropy penalties.[13, 17, 4]
@@
Ultrathin absorbers cannot be treated as step absorbers; finite absorptance reduces both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$ through reciprocity (Figure 7). Nanophotonic bounds and thickness constraints restrict the achievable broadband path-length enhancement (Figure 8). These effects motivate moderate absorber thicknesses (tens to hundreds of nm) and co-designed photonic structures rather than extreme monolayer limits for high-efficiency multijunction photovoltaics.
@@
Nonreciprocity does not improve the SQ limit for single-junction devices.[10] In multijunction architectures, however, rerouting emission pathways can reduce entropy losses and provide efficiency gain toward Landsberg-like limits.[11] Our conservative plug-in comparison (Figure 10) quantifies this efficiency gain for TMD-optimized ladders, while preserving our work’s main scope: reciprocal multijunction design and realistic constraints.
@@ 6 Conclusions
We developed an unlimited-junction detailed-balance framework tailored to transfer-printed TMD multijunction photovoltaics and used it to compute optimal bandgap ladders under unconstrained and conservative TMD bandgap windows. A realistic bandgap window produces a large-$N$ efficiency plateau, motivating $N\sim 5$ as a practical near-term target for TMD-based stacks. We incorporated luminescence thermodynamics and reciprocity to quantify ERE penalties, emission-channel entropy, and luminescent coupling. Also, we introduced a power-budget view, in which upward-emitted luminescence serves as a directly computable entropy-loss proxy, while coupling thermalization is tracked, separately. Finally, we benchmarked the remaining thermodynamic efficiency gain by applying an idealized nonreciprocal multijunction model to our reciprocally optimized ladders, showing negligible benefit for single junctions but substantial efficiency gain for multijunction stacks. These results provide quantitative design rules for approaching multicolor efficiencies with transfer-printed TMD multijunction photovoltaics.
@@
The numerical results and figures in this work are reproducible from the Python implementation of the detailed-balance, luminescent-coupling, and optimization models described in the main text and Supplement Information. The code is available at https://github.com/NEOlab-code/Multijunction-TMD-solar-cells.git. To keep the manuscript self-contained for submission, the Supplement Information includes the full derivations, parameter tables, and additional figures. The Python source code and the raw numerical data required to regenerate every figure and table are available from the authors upon reasonable request and will be deposited in a public repository upon publication.
Supplement Information
The Supplement Information is provided as a separate PDF file. The SI provides detailed derivations, parameter tables, and additional figures that support the main text:(i) Section 1 (constants, geometry, and spectral model): solar/blackbody spectral definitions and constants; (ii) Section 2 (single-junction detailed balance): full detailed-balance derivations, including the Lambert-$W$ maximum-power condition; (iii) Section 3 (unlimited-junction split-spectrum multijunction model): the dynamic-programming (DP) optimizer used to compute optimal bandgap ladders for arbitrary $N$ and its discretization/convergence checks; (iv) Section 4 (reciprocity, emission channels, and the role of mirrors): reciprocity-based voltage-loss and entropy-loss decompositions, including the role of external radiative efficiency (ERE) and optical outcoupling; (v) Section 5 (luminescent coupling chain model): the luminescent-coupling chain model, together with the definition and computation of the upward emitted luminescence power and a coupling thermalization/heat proxy; (vi) Section 6 (TMD optical response and excitonic absorptance model): a TMD material database (representative bandgaps, excitonic absorption spectra, and representative $n(\lambda)$$k(\lambda)$ dispersion conventions) and mapping of candidate stacks to the representative $N=5$ target ladder; (vii) Section 7 (nanophotonic bounds and Miller thickness constraints): nanophotonic bounds and thickness constraints relevant to ultrathin absorbers; (viii) Section 8 (nonreciprocal multijunction benchmark); (ix) Section 9 (extended bandgap-ladder tables for reproducibility); (x) Section 10 (additional figures for the representative $N=5$ stack; Figures S1 and S2); (xi) Section 11 (full ladder lists in text form).